SMC Corp. with $10m available funds is exploring uncovered interest arbitrage (UIA) opportunities under the following scenario: Spot Rate = 0.7776 ($/£) 180-day forward rate = 0.7599 ($/£) Expected Spot Rate in 180 days = 0.7752 ($/£) 180-day US interest rate = 3% 180-day UK interest rate = 1.85% The uncovered interest arbitrage potential is: Select one: a. None of the Above b. -0.31% c. 0.31% d. 1.77% e. -1.67%
Answer #1

Converting $10m into Pound = 10000000 * 0.7776 = 7776000

UK interest rate = 1.85%

Interest earned = 7776000*1.85%= Pound 143856

Total amount after 180 day = 7776000+143856 = Pound 7919856

Therefore amount in USD after 180 days = 7919856 / 0.7752 = $10216533

If amount was invested in USD then interest = 10000000 * 3% = $300000

Total amount if invested in USD = $10300000

Loss due to investment in Pound = 10300000 - 10216533 = - $83467.4923

Uncovered interest arbitage potential is = -83467.4923/10300000 = -0.008104 i.e. -0.81%

So the answer is a) None of the above